Meeting 2022 01 12
From Hall A Wiki
GMn GEM commissioning Meetings: WEDNESDAY 10:00am - 11:00am
Link to all previous meetings: GMn GEM Commissioning Meeting
Please use sbs_gems mailing list for plots/communications and HALOG entries relevant to GEMs!
Update on BB GEMs(Sean, Ezekiel, Holly, Anu)
- new UV layers having double peaking latency: 1st check if this really existed before the shutdown, also check if correlated by region (possible gas flow pockets), possibly check if correlated to gas flow, might need to check window.
- Holly following up on gas filters
Update on UVa GEM layers for GEn-RP (Anu, John, Thir)
Status & Updates
- Analyzer, SBS-offline, and SBS-replay installed on eel124gemdaq.
- Need to finalize configuration of the database and CFG files.
- One 12 slot backplane/APVs on the Inline Slack has come loose and needs to be reconnected. Involved sliding out the layer and so probably best performed once we are ready to install other layers.
- We have enough data for tracking/analysis taken with the whole setup before this issue occurred.
- Finalizing a comprehensive report on the status and remaining plans for the GEn-RP setup.
- Need some help and direction on moving forward with the SBS-Replay database and analysis.
- How should we proceed with the final layer(s) of the Left (or Right) Polarimeter?
- We have enough APVs for 4 (???) complete GEM modules - One layer with 4 installed modules is sitting in EEL125. It only needs APVs.
- What are the dates for when this gets moved in?
- finish document for Brad on status
- possible move-in date to Hall end of April/early May
- spare APVs and MPDs: how many exactly? get a box together for Paolo
- plan to tackle spare chamber?
Requests made during the meeting
- Is the MPD count in this spreadsheet (https://docs.google.com/presentation/d/1k9OGSgOSm9e2UNpA9n3k0W7nGgIyo0SK/edit#slide=id.p1) up to date? We have enough for GEn-RP?
- Re-do the inventory of APV cards. Separate out the bad APV cards, sort out which ones are fixable locally (by the electronics group or somewhere else) and the ones that might need to be sent to Italy to be fixed.
Update on INFN GEM layers for GEn-RP (Ezekiel)
- new issues with LV, but almost ready to re-test with cosmics (check for striping effect)
- can we use spare APVs from INFN and get connectors ZIF/panasonic to have more APVs for UVa chambers?
- finish document with honest assessment of current readiness
- need a plan to change out the HV divider and test
Update on the Tracking efficiency / resolution (Andrew)
Update on the High rate DAQ development (Ben, Bryan, Alex, Holly)
- Bryan implemented DAQ changes requested last week.
- Email Bryan to get an email when APVs aren't found at DAQ start.